TheSmartOne (thesmartone):

Challenge Question: Factor \(\sf a^2 + b^2=?\) Note: This isn't the Pythagorean Theorem The other similar formulas to that one, if you need it, are: \(\sf a^2-b^2 = (a+b)(a-b)\) \(\sf a^2+2ab+b^2 = (a+b)^2\) \(\sf a^2 - 2ab+b^2 = (a-b)^2\) Hint: Think outside the box ;)

OpenStudy (anonymous):

\(a^2+b^2=a^2-(-b^2)=a^2-(bi)^2=(a+bi)(a-bi)\)

OpenStudy (anonymous):

am I not supposed to post it? do I message or something?

TheSmartOne (thesmartone):

well seems like it wasn't really hard at all, great job xD

OpenStudy (usukidoll):

\[ a^2 + b^2=?\] \(\sf a^2-b^2 = (a+b)(a-b)\) \(\sf a^2+2ab+b^2 = (a+b)^2\) (I'm taking this one) \(\sf a^2 - 2ab+b^2 = (a-b)^2\) Thick outside the box? LIke this? \(\sf a^2+2ab+b^2 = (a+b)^2\) \(\sf (a+b)(a+b) = (a+b)^2\) [ using factoring] \(\sf (a+b)^2 = (a+b)^2\) [same terms, so rewrite ] I tried T_T

OpenStudy (anonymous):

this isn't rewriting but \(a^2+b^2=\frac12[(a-b)^2+(a+b)^2]\)

OpenStudy (anonymous):

err this isn't factoring but it is rewriting

TheSmartOne (thesmartone):

^

OpenStudy (usukidoll):

x( so it's like a proof... sort of.. T_T

TheSmartOne (thesmartone):

Sort of... This equation blew my mind away when my Prof told us about it xD

OpenStudy (usukidoll):

it's like.. what I'm trying to apply is that for those problems... start at the left to achieve the right or vice versa. . . .

OpenStudy (anonymous):

(a*a)+(b*b)? @TheSmartOne

TheSmartOne (thesmartone):

It's already been answered :p